Solve the system by addition.
5x-4y=-5
x-2y=-7

Answer by pvaka(23)

5x-4y=-5
5x-4y-5x=-5-5x
-4y=-5x-5
-4y/-4=-5x/-4-5/4
y=5x/4+5/4
m=5/4
y coordinate is (0,5/4) or (0,1.25)
x-2y=-7
x-2y-x=-7-x
-2y=-x-7
-2y/-2=-x/-2-7/-2
y=1x/2+7/2
m=1/2
y coordinate is (0,7/2) or (0,3.5)